2019-06-15

Saturday, 6/15 Accomplishments:

  • A/C was outside and remains outside
  • CIT-CIMS, DACOM, LGR CO & LARGE performed maintenance, software checks, plumbing
  • DIAL calibrated
  • MMS/Data performed GPS checks
  • mini DOAS ran test scans and checked communications

 
Sunday, 6/16, Plans:
Lab access only, LN2 servicing
 
Monday, 6/17, Plans:
Plane outside if DIAL needs more calibration time: tbd
0600 A/C Access, with O2 work given priority
1730 Access ends
 
Outlook:
Management is hoping the A/C is ready to fly June 24; we assume access this week, subject to O2 work priority. Adjust your travel accordingly.
We will use WhatsApp to communicate to the science team members in Palmdale. If you are here or will be soon, and don’t have the software on your phone, please install it and contact Katie Stern at 650 440 8749.  
 
ATTN: CANOE, ISAF, ROZE, WAS - Send heater system schematics to Paul Martinez/Brennan Wehr.  paul.a.martinez@nasa.gov  brennan.wehr@nasa.gov )
Send pressure system schematics to Richard Wong – even for your ground pressure systems.

2019-06-14

Dear FIREX-AQ DC-8 Experimenter,
 
Unfortunately, in the past hour the FIREX-AQ leadership was informed that the DC-8 oxygen system is broken and that the acquisition and installation of replacement parts will take a least 7 days before the DC-8 is ready to fly for the FIREX-AQ test flights.
 
Our hope is that we can perform both the FIREX-AQ Test Flights and the SARP flights during the week of June 24th.  DC-8 experimenters will likely have limited access to the DC-8 next week.  So we are recommending that DC-8 experimenters change their airline flights, hotel, and rental car accommodations such that they will be in Palmdale ready to fly a flight on Monday June 24th.
 
Our apologies for the inconvenience this is causing.  We are communicating with the ADANCEDGeo Partnership to see if/when/how to postpone Tuesday’s scheduled harassment intervention workshop.  So stay tuned for more information regarding this. The FIREX-AQ leadership will be in Palmdale next week to monitor progress, and we will be sure to give the FIREX-AQ regular updates on the status of the O2 system repair.
 
Thanks for your patience.  Best regards.
 
Barry Lefer, David Fahey, and the FIREX-AQ Leadership Team

2019-06-13

Thursday, 6/13, Accomplishments:

  • CIT-CIMS, LINeph & Master performed instrument checks, integration/software refinements, calibration, maintenance
  • DIAL prepped for calibration
  • The neph plate was installed (LARGE nephelometers, Laser Imaging neph)
  • TOGA was successfully tuned
  • DACOM completed some plumbing installs, power checks, software mods

Friday, 6/14, Plans:
Inside
0700 A/C access concurrent with O2 work. Possible intermittent power and clearing the A/C of science staff
1530 access ends
 
Outlook (changed from yesterday; each day depends on prior outcomes): 
Saturday: 0700 A/C moved outside; science access to 15:30
Sunday: CIT-CIMS (only) LN2 servicing
Monday: 0600 A/C inside, followed by science access to 1730. SARP begins.
Tuesday: address actions from Tech Brief; Preventing Harassment Training (@1400 in room 211); Potential egress training session
Wed: Shake and prof flights; potential egress training session
Thurs: Science test flight
 
ATTN: CANOE, ISAF, ROZE, WAS - Send heater system schematics to Paul Martinez/Brennan Wehr.  
(
paul.a.martinez@nasa.gov  brennan.wehr@nasa.gov)
Sign your tip over sheets. Label gas bottles.
Send pressure system schematics to Richard Wong – even for your ground pressure systems.
Empty gas bottles need to be in the empty bottle area on the back side of building 716.
Everyone should bring their own drink cup/lid for use in the lab and on flights. There is no supply of disposable cups and lids for coffee/water.
Tools taken from community A/C tool box must be returned at the end of the day.
Keep the emergency shower area in the lab free from obstructions.
RSVP to the Google survey about the June 20 Reception.

2019-06-12

Wednesday, 6/12, Accomplishments:

  • LARGE, TOGA, Mini-DOAS, CIT-CIMS, LINeph & Master performed instrument checks, integration refinements, maintenance around the O2 system work
  • DIAL prepped for calibration

Thursday, 6/13, Plans:
Still inside
0600 A/C access – continuous power until 1300.
**Please show at 0600 if you requested multiple hours of continuous power.
**O2 work concurrent; could be crowded
**After 1300, possible power interruption and clearing the A/C for O2 system work
13:00 Lab meeting (daily)
17:30 Access ends
 
Outlook (changed from yesterday): 
Friday: 07:00-15:30. Details depend on how O2 testing goes Thursday. One possibility is that the A/C is inside until Monday morning.
Saturday: 07:00-15:30 A/C access
 
ATTN: CANOE, ISAF, NOAA PTR-MS, ROZE, SP2, WAS -- Send heater system schematics to Paul Martinez/Brennan Wehr.  ( paul.a.martinez@nasa.gov  brennan.wehr@nasa.gov )
Sign your tip over sheets. Label gas bottles.
Send pressure system schematics to Richard Wong – even for your ground pressure systems.
Empty gas bottles need to be in the empty bottle area on the back side of building 716.
Everyone should bring their own drink cup/lid for use in the lab and on flights. There is no supply of disposable cups and lids for coffee/water.
Tools taken from community A/C tool box must be returned at the end of the day.
Keep the emergency shower area in the lab free from obstructions.
RSVP to the Google survey about the June 20 Reception.

2019-06-11

Tuesday, 6/11, Accomplishments:
LARGE, TOGA, Mini-DOAS, CIT CIMS, MMS, LINeph & Master performed instrument checks, integration refinements, maintenance around the O2 system work (~0930 – 1500)
DIAL prepped for calibration

Wednesday, 6/12, Plans:
06:00 A/C access (O2 work likely to start before 8; estimated 2 hrs; time your arrival accordingly)
13:00 Lab meeting (daily)
17:30 Access ends
 
Outlook (unchanged from yesterday): 
Thursday: 06:00 A/C engine run (~1 hr); then laser calibrations and A/C access until 17:30
Friday: 07:00-15:30 laser calibrations and general A/C access, outside
Saturday: 07:00-15:30 A/C access, outside, button up your instruments

ATTN: CANOE, ISAF, NOAA PTR-MS, ROZE, SP2, WAS-- Send heater system schematics to Paul Martinez/Brennan Wehr. 
Everyone should bring their own drink cup/lid for use in the lab and on flights. There is no supply of disposable cups and lids for coffee/water.
Send pressure system schematics to Richard Wong – even for your ground pressure systems.
Keep the emergency shower area in the lab free from obstructions.
RSVP to the Google survey about the June 20 Reception.

2019-06-10

Monday, 6/10, Accomplishments

  • LARGE, TOGA, Mini-DOAS, CIT CIMS, MMS & Master performed instrument checks, integration refinements, maintenance
  • DIAL prepped for calibration

Tuesday, 6/11, Plans:
06:00 – 17:30 A/C access with the following exception:
Access will occur around O2 system work, part of which requires science personnel off the A/C
13:00 Lab meeting (daily)

Outlook: 
Wednesday: 06:00 – 17:30 A/C access; again, possible interruption for O2 system work
Thursday: 06:00 A/C engine run (~1 hr); then laser calibrations and A/C access until 17:30
Friday: 07:00-15:30 laser calibrations and general A/C access, outside
Saturday: 07:00-15:30 A/C access, outside, button up your instruments

Very high temps expected Tuesday (100°F +).
 

2019-06-08

Saturday, 6/8, Accomplishments:
Instrument integration tasks; occasionally interupted by O2 system work

Sunday 6/9, Plans:
No A/C access
Lab access

2019-06-07

Friday, 6/7, Accomplishments:
Instrument integration
MASTER loaded in aft cargo pit
A/C O2 system checks

Saturday, 6/8, Plans:
08:00 - 09:30    A/C Access
09:30 - 11:00 (an estimate of completion time) A/C O2 system work, no A/C access
After the completion of the O2 system work there will be A/C access until 15:30

Send pressure system schematics to Richard Wong

Send heater system schematics to Paul Martinez/Brennan Wehr

2019-06-06

Thursday, 6/6, Accomplishments:
OPALS team performed probe fit check
ACES/IWAS rack installed
TOGA continued working on sampling lines between inlet and instrument
DIAL team made good progress
mini-DOAS team performed in-lab calibrations
NO LIF instrument removed from rack; to be reinstalled on the 18th

Friday, 6/7, Plans:
06:00 A/C Access
MASTER to load aft pit
Instrument integration tasks
13:00 Lab meeting
17:30 A/C access ends

Send pressure system schematics to Richard Wong

Send heater system schematics to Paul Martinez/Brennan Wehr

2019-06-05

Wednesday, 6/5, Accomplishments:
OPALS team performed in lab installation prep
mini DOAS team performed in lab calibrations
ACES and IWAS testing in lab, waiting for the install window
TOGA installed
ISAF/ROZE/CANOE and SO2/NO-LIF installed
CIT-CIMS rack and pump plate installed
DIAL completed nadir alignments

Thursday, 6/6, Plans:
06:00 A/C Access
IWAS/ACES install
Instrument installations and integration
13:00 Lab meeting
17:30 A/C access ends

Tools taken from community A/C tool box must be returned at the end of the day
Label gas bottles
Empty gas bottles need to be in the empty bottle area on the back side of building 716
Send pressure system schematics to Richard Wong
Send heater system schematics to Paul Martinez/Brennan Wehr

Pages

Subscribe to Daily Schedule